OpenStudy (anonymous):

Add the numbers starting from 1 up to 100.then what do u get. Ex.1+2+3+4+5+6.................100, what do u get ???????

OpenStudy (freckles):

use the following to find the sum: \[1+2+3+\cdots +n=\frac{n(n+1)}{2}\]

geerky42 (geerky42):

Let \(S = 1+2+3+4+\cdots+100\) \[~~1~~+~2~+~3~+~4~+\cdots+100\\~\\100+99+98+97+\cdots+~~1\] \[2S = \underbrace{101+101+101+\cdots+101}_{\Large100~times} = 100\cdot101\] So \(S = \dfrac{100\cdot101}{2} = \cdots~?\)

OpenStudy (freckles):

or in general \[S_n=1+2+3+\cdots+(n-2) +(n-1)+n \\S_n=n+(n-1)+(n-2)+\cdots+3+2+1\\ S_n+S_n \\ =(n+1)+(n-1+2)+(n-2+3)+\cdots +(n-2+3)+(n-1+2)+(n+1) \\ 2 S_n=(n+1)+(n+1)+(n+1) \cdots +(n+1)+(n+1)+(n+1) \\ 2S_n=n(n+1) \\ S_n=\frac{n(n+1)}{2}\] \[1+2+3+\cdots +n =\frac{n(n+1)}{2} \]